# Building Math Agents with Multi-Turn Iterative Preference Learning

### v1

@@ Abstract
Recent studies have shown that large language models’ (LLMs) mathematical problem-solving capabilities can be enhanced by integrating external tools, such as code interpreters, and employing multi-turn Chain-of-Thought (CoT) reasoning. While current methods focus on synthetic data generation and Supervised Fine-Tuning (SFT), this paper studies the complementary direct preference learning approach to further improve model performance. However, existing direct preference learning algorithms are originally designed for the single-turn chat task, and do not fully address the complexities of multi-turn reasoning and external tool integration required for tool-integrated mathematical reasoning tasks. To fill in this gap, we introduce a multi-turn direct preference learning framework, tailored for this context, that leverages feedback from code interpreters and optimizes trajectory-level preferences. This framework includes multi-turn DPO and multi-turn KTO as specific implementations. The effectiveness of our framework is validated through training of various language models using an augmented prompt set from the GSM8K and MATH datasets. Our results demonstrate substantial improvements: a supervised fine-tuned Gemma-1.1-it-7B model’s performance increased from 77.5% to 83.9% on GSM8K and from 46.1% to 51.2% on MATH. Similarly, a Gemma-2-it-9B model improved from 84.1% to 86.3% on GSM8K and from 51.0% to 54.5% on MATH.
@@ 1 Introduction
Built on strong SFT models, Reinforcement Learning from Human Feedback (RLHF) has proven to be a key technique to elicit LLMs’ knowledge during the post-training stage and has become a standard practice in the LLM training pipeline (Bai et al., 2022; Ouyang et al., 2022; Touvron et al., 2023; Team et al., 2023). Broadly speaking, the RLHF learning paradigm, which was originally designed for aligning large language models (LLMs) with human values and preferences (Bai et al., 2022; Ouyang et al., 2022), is distinct from SFT as it learns from relative feedback (Christiano et al., 2017; Ziegler et al., 2019). It has notably enhanced the capabilities of models like ChatGPT, Claude, and Gemini, enabling them to generate responses that are more helpful, harmless, and honest (Bai et al., 2022). Inspired by RLHF’s success in general chat applications, in this paper, we explore RLHF for improving LLMs’ mathematical problem-solving abilities when equipped with external tools. In particular, since deep RL methods (e.g., the proximal policy optimization, PPO algorithm (Schulman et al., 2017)) are often sample inefficient and unstable (Choshen et al., 2019), our goal is to derive direct preference learning algorithms that directly learn from the preference dataset (Zhao et al., 2023; Rafailov et al., 2023; Azar et al., 2023).
Contribution.
We begin by formulating the learning process as a Markov decision process (MDP), distinct from the contextual bandit approach typically used in RLHF for making general chatbots without external environment interactions (Xiong et al., ; Rafailov et al., 2023). Then, we derive the optimality condition of the optimization problem and develop multi-turn direct alignment algorithms (M-DPO and M-KTO) that incorporate external messages, where the primary modification is to mask out irrelevant tokens during training. Furthermore, we extend our approach to its online iterative variants, which recent works demonstrated to be promising (Xiong et al., ; Guo et al., 2024b). Finally, we evaluate our approach through case studies using augmented training sets from MATH and GSM8K benchmarks, employing various base models such as Gemma (Team et al., 2024), CodeGemma (Team, 2024), and Mistral (Jiang et al., 2023). For instance, the performance of a supervised fine-tuned Gemma-1.1-it-7B model increased from 77.5% to 83.9% on GSM8K and from 46.1% to 51.2% on MATH. Similarly, a Gemma-2-it-9B model improved from 84.1% to 86.3% on GSM8K and from 51.0% to 54.5% on MATH. These empirical results indicate a significant improvement in performance over standard SFT models, demonstrating the potential of RLHF in complex reasoning task. We also provide a comprehensive recipe for the practical implementation of our online iterative multi-turn methods, and make our models, datasets, and code publicly available for further research and development.
@@ 1.1 Problem Formulation
We denote prompt as $x\in\mathcal{X}$ and assume that the interactions run for up to $H$ rounds. At the first step, a prompt $x$ is sampled from some distribution $d_{0}$ as the initial state $s_{1}=x$ (We use the terminology “state” instead of “context” because we are concerning about an MDP instead of a contextual bandit here). Then, at each step $h\in[H]$,
@@
See Figure 1 for an example. The framework presented here is a Markov decision process (MDP), which offers a distinct approach from the contextual bandit model discussed in Xiong et al. . Formally, we define the following MDP.
@@
Figure 1: An example of multi-turn mathematical reasoning with Python interpreter. The action is in a ReAct style (Yao et al., 2022) where it consists of a reasoning step $f_{h}$ and an executation step $e_{h}$.
@@ 2.1 Planning with a Fixed Model: Optimality Condition
Moving from the single-step to multi-turn scenario, we first show that we are still concerning about the Gibbs distribution, but in a dynamic programming manner. The results are essentially from the study of entropy-regularized MDPs (Williams and Peng, 1991; Ziebart, 2010).
@@ 2.2 Planning with a Fixed Model: Practical Algorithm
Here, term $(A)$ is the counterpart of $\eta\log\frac{\pi(a_{1}|s_{1})}{\pi_{\mathrm{ref}}(a_{1}|s_{1})}$ in the single-step DPO derivation and term $(B)$ will be cancelled if we consider the reward difference of two trajectories with the same prompt $s_{1}=x$. Unfortunately, in practice, term $(C)$ is typically not feasible to directly compute. Especially, some simple math with the Chebyshev’s Inequality leads to that with probability at least $0.9$,
@@ 2.3 Learning with Online Iterative Training
We present a general online iterative algorithmic framework in Algorithm 1. This framework is termed as Online Iterative Multi-turn Gibbs Sampling from Human Feedback (M-GSHF) to highlight the online iterative training process and the optimal condition derived in (7) that the optimal policy is a layer-wise Gibbs distribution, which generalizes the bandit formulation in Xiong et al. . Specifically, starting from $\pi_{0}$, at each iteration, we first collect a pair of trajectories by the current policy pair, where the preference signal is also revealed according to Definition 1. Then, we update our policy pair given the data collected so far and the next iteration begins. We now discuss some features of the framework as follows.
@@
12:     if KL-regularized target in (14then
@@
14:     else if Non-regularized target in (15then
@@ 3.2 Main Results
Table 1: Main results of different methods on the test sets of GSM8K and MATH. The SFT training with external tool is based on (a subset of) Open-MathInstruct so the results are generally comparable to the previous SFT models. $\dagger$: the model also serves as the starting checkpoint of other methods except for prompting and CoT without tool use. All the models are allowed to use code interpreter except for the CoT without tool use. The results of the CoT methods are borrowed from the technical reports (Toshniwal et al., 2024; Gou et al., 2023b). The gains relative to the SFT starting checkpoint is marked by  $\uparrow$.
@@
We also observe that the iterative M-DPO and M-KTO surpass other existing RLHF baselines. First, they consistently and significantly outperform the RAFT algorithm across all four base models, which is known to be a robust and competitive baseline in the literature (Dong et al., 2023; Yuan et al., 2023a). This is because the RAFT algorithm only utilizes the positive signal by imitating the correct trajectories, while the DPO-based and KTO-based algorithms further leverage the negative signal from those incorrect trajectories. We note that the SFT stage in our pipeline can also be viewed as an application of RAFT, an idea that further dates back to expert iteration (Anthony et al., 2017). Consequently, our results should be interpreted to be that on the top of the first stage of SFT, algorithms with negative signal are more sample efficient. Moreover, while the online iterative single-turn DPO (KTO) (Xiong et al., ; Xu et al., 2023) also gives a boost performance, it is generally worse than the multi-turn version. This suggests that learning to predict the off-policy external messages returned by the code interpreter usually has a negative impact on the reasoning ability improvement. Essentially, this corresponds to the fact that when deriving the optimality condition of the KL-regularized optimization problem, we are not allowed to optimize the external messages. Meanwhile, we present a representative example we encounter in Figure 3, where LLMs generate poorly constructed code resulting in anomalous and lengthy external messages. Forcing LLMs to learn to predict these messages can significantly hurt the model’s reasoning abilities.
Figure 3: An example of external messages returned by the Python interpreter. The model writes down a bad python code leading to an anomalous and lengthy error message.
@@ Appendix B Implementation Detail
Following Gou et al. (2023b); Toshniwal et al. (2024), the LLM agent is allowed to call the python interpreter when it decodes a python code starting with ```python and ending with ```. For each step $h$, to generate the observation $o_{h}$, we leverage the python package IPython, and run all the codes in the history one by one and treat each code snippet as a Jupyter cell. We only return the standard output or the error message from the last snippet. When there exists some bug in the code, we only return the error message which is typically less than 20 tokens as in Toshniwal et al. (2024). We notice that some works (e.g. Shao et al. (2024)) also returns the first and the last 50 tokens of the trackback information.
@@
We do not tune the prompt though we do observe that the prompt engineering can further improve the performance. For all the experiments, we simply adopt the chat template of the models as in Figure 1.

### v2

@@ Abstract
Recent studies have demonstrated the potential to enhance the mathematical problem-solving capabilities of large language models (LLMs) by integrating external tools such as code interpreters and employing multi-turn Chain-of-Thought (CoT) reasoning. While existing approaches primarily focus on synthetic data generation and Supervised Fine-Tuning (SFT), this paper explores complementary preference learning to further improve model performance. However, existing direct preference learning algorithms are originally designed for the single-turn chat task, and do not fully address the complexities of multi-turn reasoning and external tool integration required for tool-integrated mathematical reasoning tasks. To fill in this gap, we introduce a multi-turn online iterative direct preference learning framework tailored to this unique context, which incorporates feedback from code interpreters and optimizes trajectory-level preferences. The effectiveness of our framework is validated through training of various language models using an augmented prompt set derived from GSM8K and MATH datasets. Our results show significant improvements even with only final result checking: for instance, the performance of a supervised fine-tuned Gemma-1.1-it-7B model increased from 77.5% to 83.9% on GSM8K and from 46.1% to 51.2% on MATH. Similarly, a Gemma-2-it-9B model improved from 84.1% to 86.3% on GSM8K and from 51.0% to 54.5% on MATH.
@@ 1 Introduction
Building on strong SFT models, Reinforcement Learning from Human Feedback (RLHF) has proven to be a key technique to elicit LLMs’ knowledge during the post-training stage and has become a standard practice in the LLM training pipeline (Bai et al., 2022; Ouyang et al., 2022; Touvron et al., 2023; Team et al., 2023). Broadly speaking, the RLHF learning paradigm, which was originally designed for aligning large language models (LLMs) with human values and preferences (Bai et al., 2022; Ouyang et al., 2022), is distinct from SFT as it learns from relative feedback (Christiano et al., 2017; Ziegler et al., 2019). It has notably enhanced the capabilities of models like ChatGPT, Claude, and Gemini, enabling them to generate responses that are more helpful, harmless, and honest (Bai et al., 2022). Inspired by RLHF’s success in general chat applications, in this paper, we explore RLHF for improving LLMs’ mathematical problem-solving abilities when equipped with external tools.
In particular, since deep RL methods (e.g., the proximal policy optimization, PPO algorithm (Schulman et al., 2017)) are often sample inefficient and unstable (Choshen et al., 2019), our goal is to derive direct preference learning algorithms that directly learn from the preference dataset (Zhao et al., 2023; Rafailov et al., 2023; Azar et al., 2023). We begin by formulating the learning process as a Markov decision process (MDP), distinct from the contextual bandit approach typically used in RLHF for making general chatbots without external environment interactions (Xiong et al., ; Rafailov et al., 2023). Then, we derive the optimality condition of the optimization problem and develop multi-turn variants of direct preference learning algorithms that incorporate external messages, where the primary modification is to mask out irrelevant tokens during training. Furthermore, we extend our approach to its online iterative variants, which recent works demonstrated to be promising (Xiong et al., ; Guo et al., 2024b).
We evaluate our approach through case studies using augmented training sets from MATH and GSM8K benchmarks, employing various base models such as Gemma (Team et al., 2024), CodeGemma (Team, 2024), and Mistral (Jiang et al., 2023). For instance, the performance of a supervised fine-tuned Gemma-1.1-it-7B model increased from 77.5% to 83.9% on GSM8K and from 46.1% to 51.2% on MATH. Similarly, a Gemma-2-it-9B model improved from 84.1% to 86.3% on GSM8K and from 51.0% to 54.5% on MATH. These empirical results indicate a significant improvement in performance over standard SFT models, demonstrating the potential of RLHF in complex reasoning task. We also provide a comprehensive recipe for the practical implementation of our online iterative multi-turn methods, and make our models, datasets, and code publicly available for further research and development.
@@ 1.1 Problem Formulation
We denote prompt as $x\in\mathcal{X}$ and assume that the interactions run for up to $H$ rounds. At the first step, a prompt $x$ is sampled from some distribution $d_{0}$ as the initial state $s_{1}$ (We use the terminology “state” instead of “context” because we are concerning about an MDP instead of a contextual bandit here). Then, at each step $h\in[H]$,
@@
See Table LABEL:tab:multi_turn_math for an example. The framework presented here is a Markov decision process (MDP), which offers a distinct approach from the contextual bandit model discussed in Xiong et al. . Formally, we define the following MDP.
@@ 2.1 Planning with a Fixed Model: Optimality Condition
Moving from the single-step to multi-turn scenario, we first show that we are still concerning about the Gibbs distribution, but in a dynamic programming manner. We summarize the results into the following proposition.
Proposition 1.
We can recursively define the following optimal value functions and optimal policies for a KL-regularized MDP with horizon $H$ and external observation $o_{h}$. For Q value, we have
$Q_{\mathcal{M},h}(s_{h},a_{h})=\begin{cases}&u(s_{H},a_{H}),\qquad\text{ if }h=H,\\ &\mathbb{E}_{o_{h}\sim\mathbb{P}_{h}(\cdot|s_{h},a_{h})}[V_{\mathcal{M},h+1}(s_{h+1})],\qquad\text{ if }h\leq H-1.\end{cases}$ (6)
Also, for all $h\in[H]$, we have:
$V_{\mathcal{M},h}(s_{h})=\eta\log\underbrace{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbb{E}_{a_{h}\sim\pi_{\mathrm{ref},h}(\cdot\mid s_{h})}}\exp\Big{(}\frac{Q_{\mathcal{M},h}(s_{h},a_{h})}{\eta}\Big{)}}_{=:Z_{h}(s_{h})},$
$\pi_{\mathcal{M},h}(a_{h}\mid s_{h})=\frac{\pi_{\mathrm{ref},h}(a_{h}\mid s_{h})}{Z_{h}(s_{h})}\cdot\exp\Big{(}\frac{Q_{\mathcal{M},h}(s_{h},a_{h})}{\eta}\Big{)}.$ (7)
We have a few interesting observations that may be of independent interests.
  1. 1.
    The optimal value function is characterized by the expectation with respect to the initial reference policy due to the additional KL constraint.
  2. 2.
    For a fixed step $h$ and state-action pair $(s_{h},a_{h})$, we can treat the future as a bandit (with only one step), then, we have $Q_{\mathcal{M},h}(s_{h},a_{h})=\mathbb{E}_{z}u(s_{h},a_{h},z)$, where $z$ is a completion staring from $(s_{h},a_{h})$. One can use the Monte-Carlo estimation to estimate this value by multiple roll-outs. We notice that the non-regularized version of this process, is commonly referred to as the process-supervised reward (PRM) in the literature (Wang et al., 2023a). In other words, the PRM constructed in Wang et al. (2023a) is essentially a Q learning process.
The results are essentially from the study of entropy-regularized MDPs (Williams and Peng, 1991; Ziebart, 2010).
@@ 2.2 Planning with a Fixed Model: Practical Algorithm
Here, term $(A)$ is the counterpart of $\eta\log\frac{\pi(a_{1}|s_{1})}{\pi_{\mathrm{ref}}(a_{1}|s_{1})}$ in the single-step DPO derivation and term $(B)$ will be cancelled if we consider the reward difference of two trajectories with the same prompt $s_{1}=x$. Unfortunately, in practice, term $(C)$ is typically not feasible to directly compute. Especially, some simple math leads to that with probability at least $0.9$,
@@ 2.3 Learning with Online Iterative Training
We present a general online iterative algorithmic framework in Algorithm 1. Specifically, starting from $\pi_{0}$, at each iteration, we first collect a pair of trajectories by the current policy pair, where the preference signal is also revealed according to Definition 1. Then, we update our policy pair given the data collected so far and the next iteration begins. We now discuss some features of the framework as follows.
@@
12:     if KL-regularized target in (16then
@@
14:     else if Non-regularized target in (17then
@@ 3.2 Main Results
Table 1: Main results of different methods on the test sets of GSM8K and MATH. The SFT training with external tool is based on (a subset of) Open-MathInstruct so the results are generally comparable to the previous SFT models. $\dagger$: the model also serves as the starting checkpoint of other methods except for prompting and CoT without tool use. All the models are allowed to use code interpreter except for the CoT without tool use. The results of the CoT methods are borrowed from the technical reports (Toshniwal et al., 2024; Gou et al., 2023b). The gains relative to the SFT starting checkpoint are marked by  $\uparrow$.
@@
We also observe that the iterative M-DPO and M-KTO surpass other existing RLHF baselines. First, they consistently and significantly outperform the RAFT algorithm across all four base models, which is known to be a robust and competitive baseline in the literature (Dong et al., 2023; Yuan et al., 2023a). This is because the RAFT algorithm only utilizes the positive signal by imitating the correct trajectories, while the DPO-based and KTO-based algorithms further leverage the negative signal from those incorrect trajectories. We note that the SFT stage in our pipeline can also be viewed as an application of RAFT, an idea that further dates back to expert iteration (Anthony et al., 2017). Consequently, our results should be interpreted to be that on the top of the first stage of SFT, algorithms with negative signal are more sample efficient. Moreover, while the online iterative single-turn DPO (KTO) (Xiong et al., ; Xu et al., 2023) also gives a boost performance, it is generally worse than the multi-turn version. This suggests that learning to predict the off-policy external messages returned by the code interpreter usually has a negative impact on the reasoning ability improvement. Essentially, this corresponds to the fact that when deriving the optimality condition of the KL-regularized optimization problem, we are not allowed to optimize the external messages. Meanwhile, we present a representative example we encounter in Table LABEL:tab:example_observation, where LLMs generate poorly constructed code resulting in anomalous and lengthy external messages. Forcing LLMs to learn to predict these messages can significantly hurt the model’s reasoning abilities.
@@ 4 Conclusion, Limitation, and Future Research Direction
Acknowledgements
Wei Xiong and Tong Zhang are partially supported by an NSF IIS grant No. 2416897
@@ Appendix B Implementation Detail
Following Gou et al. (2023b); Toshniwal et al. (2024), the LLM agent is allowed to call the python interpreter when it decodes a python code starting with “‘python and ending with “‘. For each step $h$, to generate the observation $o_{h}$, we leverage the python package IPython, and run all the codes in the history one by one and treat each code snippet as a Jupyter cell. We only return the standard output or the error message from the last snippet. When there exists some bug in the code, we only return the error message which is typically less than 20 tokens as in Toshniwal et al. (2024). We notice that some works (e.g. Shao et al. (2024)) also returns the first and the last 50 tokens of the trackback information.
@@
We do not tune the prompt though we do observe that the prompt engineering can further improve the performance. For all the experiments, we simply adopt the chat template of the models as in Table LABEL:tab:multi_turn_math.